# DiffFuSR: Super-Resolution of all Sentinel-2 Multispectral Bands using Diffusion Models

### v1

@@ V Discussion
(a) High-Resolution (Reference)
(b) DiffFuSR (NAIP Harmonized)
(c) OpenSR
(d) Low-Resolution (Input)
(e) LDM_Baseline
(f) Satlas
(g) SuperImage
Figure 7: Visual comparison of SR methods on OpenSR-test  [15] test image. The top row highlights the strongest models: HR reference, our model trained on harmonized NAIP, and OpenSR. The bottom row shows the input LR image and remaining baselines. Among all methods, DiffFuSR produces the sharpest reconstruction with fewer artifacts, while OpenSR is a strong baseline. LDM_Baseline and Satlas produce visibly degraded outputs, and SuperImage tends to oversmooth fine structures.
We presented DiffFuSR, a modular pipeline to super-resolve all 12 spectral bands of Sentinel-2 imagery to a 2.5 m GSD, leveraging a diffusion-based model for RGB SR and a learned network for multispectral fusion. The evaluation of the proposed RGB SR model against several contemporary baselines on the OpenSR-test benchmark demonstrated its capabilities in quantitative terms. However, the qualitative results need to be paired with quantitative results to accurately gauge the true performance. Looking at one of the columns of results in quantitative Tables can be misleading, e.g., Satlas and LDM_Baseline seem to be doing the best in terms of Omission metric in Table III. But when looking at the rest of the values and also the qualitative results in Fig. 7, the SR outputs contain severe artefacts.
When trained with harmonized NAIP data, the proposed diffusion model consistently demonstrated superior or highly competitive performance across multiple quantitative metrics, including Normalized Difference, CLIP distance, and LPIPS (Tables I - III). For instance, it achieved the lowest reflectance error (0.0024) and highest spectral fidelity (1.1103) in Normalized Difference metrics, and the best spectral score (1.6739) with CLIP distance. It also recorded a competitive hallucination score (0.1149) using LPIPS, alongside one of the best improvement scores (0.5383) with the same metric, indicating an enhanced ability to reconstruct rich, meaningful high-frequency content with minimal introduction of spurious details. Qualitatively, our model, particularly the harmonized NAIP variant, produced sharper reconstructions with fewer artifacts compared to strong baselines like OpenSR, and markedly better results than LDM_Baseline, Satlas, or Superimage, which exhibited issues such as instability, poor spectral integrity, or over-smoothing, respectively (Fig. 7).
@@
For the subsequent enhancement of all 12 Sentinel-2 bands, the learned multispectral fusion network was evaluated against classical GS pansharpening. The results were definitive: the neural network (NN)-based fusion outperformed GS across all considered metrics (R2̂, Cross-Correlation, SSIM, PSNR, MSE, and ERGAS) and for all RGB SR inputs (Tab. IV). For example, when using the harmonized NAIP SR RGB, ERGAS values dropped from over 2900 for GS to under 95 for our NN approach, and PSNR increased from approximately 22 dB to over 32 dB. Qualitative comparisons (Fig. 6) further confirmed that our fusion model better preserves semantic structures and avoids the spectral distortions common in GS fusion, especially for the 20 m and 60 m bands.
@@
A critical factor in the success of the RGB SR stage was the spectral harmonization and degradation modeling applied to the proxy HR datasets (NAIP and WorldStrat) to simulate Sentinel-2 characteristics. The superior performance of models trained on harmonized NAIP data, particularly for vegetated scenes (Fig. 4), underscores the importance of these pre-processing steps. The blind degradation modeling, facilitated by a contrastive degradation encoder, is a significant contribution, enabling the SR model to adapt to unseen degradation conditions in real Sentinel-2 imagery without explicit kernel information.
@@
The choice of HR training data also influenced performance, with NAIP-trained models excelling in natural, vegetated environments, while WorldStrat-trained models demonstrated an advantage in structured, urban settings (Figs. 4 and 5). This suggests that while harmonization is effective, the inherent spatial and radiometric properties of the proxy data can still subtly influence model specialization, highlighting the value of diverse training datasets for developing globally applicable models.
The effect of harmonization on NAIP data is most prominent on reflectance and spectral metrics. Other metrics are not affected to the same extent. This is in accordance to our expectation as harmonization brings the spectral content of NAIP closer to Sentinel-2. It is also interesting to note that metrics like hallucination are also slightly reduced, giving interesting insight into the effect of harmonization. On the other hand, the omission metric is actually better in the unharmonized case. This behavior is interesting because it indicates that the SR images have missing details that were present in the HR image.
@@
The learned fusion network, trained using only native Sentinel-2 data via a Wald protocol simulation, proved highly effective. Its ability to utilize the 2.5 m super-resolved RGB output from the diffusion model and to upscale the 10 m, 20 m, and 60 m bands while preserving spectral integrity is a core achievement. The scale-aware design, with dedicated fusion modules and tailored inputs for each Sentinel-2 resolution group (10 m, 20 m, 60 m bands), contributes to the robustness and high fidelity of the final super-resolved Sentinel-2 12-band product. This modularity, both within the fusion stage and in the overall two-stage pipeline, offers practical benefits for future refinement and adaptation. There is room for improvement in the design choices in the Fusion module architecture, which we leave for future work.
VI Conclusions
@@
In this work, we proposed and demonstrated DiffFuSR, a pipeline to super-resolve all 12 bands of Sentinel-2. The main contribution of the work was the combination of strengths of both classical and state-of-the-art deep learning methods, and ensuring best practices for pre-processing and harmonizing the data. The work has demonstrated that learning state-of-the-art Sentinel-2 SR models from other data modalities, e.g., NAIP, is not only possible but one of the best ways due to the lack of paired data. Converting the NAIP data to look like Sentinel-2 is the most critical factor in the pipeline. Small details like the correct kernel size and the type of blur kernel are crucial factors for an accurate degradation model to simulate Sentinel-2 images. By learning from NAIP data with the correct degradation model, we were able to train a diffusion model to super-resolve the RGB bands of Sentinel-2. To super-resolve all 12 bands, the super-resolved RGB image was used to guide the enhancement of the remaining bands in the proposed fusion module.
However, some limitations and areas for future development should be acknowledged. The computational requirements for training and, more critically, for inference with diffusion models are substantial due to their iterative sampling nature. For widespread operational adoption, research into model compression, knowledge distillation, or more efficient sampling strategies specifically for SR tasks will be crucial. While the training datasets (SEN2NAIP, WorldStrat) are extensive, developing models that exhibit even greater robustness across the full diversity of global landscapes and atmospheric conditions will likely benefit from even larger and more varied training corpora, potentially incorporating advanced data augmentation or domain generalization techniques. The reliance on proxy data, even with careful harmonization, may also leave residual domain gap effects that warrant continued research into cross-domain adaptation.
Future work could extend this framework to incorporate temporal information from Sentinel-2 time series, potentially improving SR quality and enabling consistent 2.5 m analysis-ready data stacks. Adapting the degradation models and fusion strategies for other multispectral satellite sensors could broaden the applicability of this promising pipeline architecture.

### v2

@@ V Discussion
HR (Reference)
DiffFuSR (NAIP Harmonized)
OpenSR
LDM_Baseline
Satlas
SuperImage
Figure 9: Visual comparison of SR methods on an OpenSR-test [opensrtest] image. The top row highlights the strongest models: HR reference, our model trained on harmonized NAIP, and OpenSR. The bottom row shows the remaining baselines. Among all methods, DiffFuSR produces the sharpest reconstruction with fewer artifacts, while OpenSR remains a strong baseline. LDM_Baseline and Satlas show degraded outputs, and SuperImage tends to oversmooth fine structures.
We note that the qualitative results need to be paired with quantitative results to accurately gauge the true performance. Looking at one of the columns of results in quantitative Tables can be misleading, e.g., Satlas and LDM_Baseline seem to be doing the best in terms of Omission metric in Table III. But when looking at the rest of the values and also the qualitative results in Fig. 9, the SR outputs contain severe artefacts.
The strong performance of the harmonized NAIP model across all metric categories (Tables I - III) highlights the critical role of spectral harmonization. By aligning the source domain (NAIP) with the target (Sentinel-2), the model achieves superior reflectance and spectral fidelity, indicating it learns a more accurate representation of Sentinel-2’s characteristics. Qualitatively, particularly the harmonized NAIP variant, produced sharper reconstructions with fewer artifacts compared to strong baselines like OpenSR, and markedly better results than LDM_Baseline, Satlas, or Superimage, which exhibited issues such as instability, poor spectral integrity, or over-smoothing, respectively (Fig. 9).
@@
However, the superior performance of the GLP baseline when matching low-resolution images does not translate into qualitative results. As shown in Fig. 7, the GLP baseline produces very blurry outputs, whereas our proposed method generates sharp results with well-preserved details while remaining the most accurate in terms of quantitative metrics.
@@
A critical factor in the success of the RGB SR stage was the spectral harmonization and degradation modeling applied to the proxy HR datasets (NAIP and WorldStrat) to simulate Sentinel-2 characteristics. The blind degradation modeling, facilitated by a contrastive degradation encoder, is a significant contribution, enabling the SR model to adapt to unseen degradation conditions in real Sentinel-2 imagery without explicit kernel information.
@@
This suggests that while harmonization is effective, the inherent spatial and radiometric properties of the proxy data can still subtly influence model specialization, highlighting the value of diverse training datasets for developing globally applicable models.
@@
The learned fusion network, trained using only native Sentinel-2 data via a Wald protocol simulation, proved highly effective. Its ability to utilize the 2.5 m super-resolved RGB output from the diffusion model and to upscale the 10 m, 20 m, and 60 m bands while preserving spectral integrity is a core achievement. The scale-aware design, with dedicated fusion modules and tailored inputs for each Sentinel-2 resolution group (10 m, 20 m, 60 m bands), contributes to the robustness and high fidelity of the final super-resolved Sentinel-2 12-band product. This modularity, both within the fusion stage and in the overall two-stage pipeline, offers practical benefits for future refinement and adaptation. There is room for improvement in the design choices in the Fusion module architecture, to match the quantitative performance of the GLP baseline, which we leave for future work.
V-A Implications of the Evaluation Strategy
A key strength of this study’s SR evaluation is the use of the OpenSR-test benchmark. By assessing performance across three different distance metrics (Normalized Difference, LPIPS, and CLIP), we can move beyond simple pixel error. This provides a more holistic validation, capturing physical accuracy (reflectance), perceptual realism (LPIPS), and semantic consistency (CLIP), which is critical for gauging the true utility of generative models.
It is crucial to properly interpret the quantitative results for the multispectral fusion (Tab. IV). In the absence of high-resolution ground truth for all 12 Sentinel-2 bands, performance was assessed indirectly by downsampling the super-resolved outputs back to their native resolution and comparing them there. This ”no-reference” validation, based on the Wald protocol, primarily measures fidelity to the original low-resolution signal, not absolute accuracy at the 2.5 m target resolution. This limitation explains the apparent discrepancy noted in our results: the classical GLP method excelled in the quantitative metrics (by producing a blurry result very similar to the downsampled input) but was clearly inferior in qualitative assessment (Fig. 7), where it produces very blurry outputs. Our proposed GLP-NN, therefore, strikes a better balance, enhancing spatial detail while remaining acceptably consistent with the original spectral information.
VI Limitations and Future Work
While DiffFuSR advances Sentinel-2 super-resolution, several limitations remain that can be addressed in future work:
  • Proxy supervision and domain gap: Training the RGB SR model on NAIP and WorldStrat data with statistics-based harmonization cannot fully remove sensor and illumination differences. Residual domain gaps may bias colors and textures in super-resolved outputs. Future research could explore large-scale datasets of harmonized cross-sensor imagery, such as extended SEN2NAIP [aybar2024sen2naip] variants with more temporal diversity, which could help model real-world variations more effectively.
  • Guidance bias in fusion: Using SR-RGB as a spatial prior can transfer RGB-specific artifacts or impose structures not present in non-RGB bands (e.g., SWIR), risking subtle spectral distortions despite our safeguards. Future work may investigate fusion models with adaptive spectral weighting to dynamically regulate RGB influence per band.
  • Evaluation constraints: In the absence of HR Sentinel-2 ground truth, we rely on OpenSR-style proxies and Wald-based no-reference validation; these may not perfectly reflect absolute radiometric fidelity in the wild. Future evaluation can come up with suitable downstream applications that demonstrate the utility of the super-resolved bands
VII Conclusions
@@
In this work, we proposed and demonstrated DiffFuSR, a pipeline to super-resolve all 12 bands of Sentinel-2. The main contribution of the work was the combination of strengths of both classical and state-of-the-art deep learning methods, and ensuring best practices for pre-processing and harmonizing the data. The work has demonstrated that learning state-of-the-art Sentinel-2 SR models from other data modalities, e.g., NAIP, is not only possible but one of the best ways due to the lack of paired data. Converting the NAIP data to look like Sentinel-2 is the most critical factor in the pipeline. Small details like the correct kernel size and the type of blur kernel are crucial factors for an accurate degradation model to simulate Sentinel-2 images. By embedding the principles of modulation transfer function filtering and regression-based detail injection into a learnable architecture, the model inherits the interpretability and stability of traditional approaches while benefiting from the adaptability of deep learning.
Beyond the technical contributions, DiffFuSR represents a practical step toward bridging the gap between open-access data and costly high-resolution commercial imagery. By providing a reliable method to generate a 2.5 m, 12-band Sentinel-2 product, this work can unlock new capabilities for a wide range of applications, such as detailed land cover mapping, agricultural monitoring, and urban analysis, for researchers who rely solely on the public Copernicus data stream. Ultimately, this work demonstrates that a modular pipeline, thoughtfully combining generative models with classical principles and rigorous data handling, is key to enhancing the value of foundational EO datasets like Sentinel-2.
@@ Acknowledgments
VIII Supplementary Materials
Fig. 10 and  11 show results that are supplementary to the ones in the main paper. They evaluate the proposed learnable fusion model. Here, we only use the harmonised NAIP-based SR model to create the RGB super-resolved image that is further used for fusion. The columns represent the Low Resolution input, Generalized laplacian pyramid (GLP) baseline method [vivone2018full], our GLP inspired neural network, intensity-hue-saturation (IHS) transform, and Fusion network baseline [deng2022machine] and Gram Schmidt [dalla2015global] method applied to 60 m bands and visualized in RGB, where the red and green channels use the Water Vapour band (B10) and the blue channel uses the Coastal Aerosol band (B1). Compared to the baselines, which often distort the structure and exaggerate spectral contrast, our GLP inspired fusion networks output is visually more consistent with the low-resolution signal while enhancing spatial detail.
Low Resolution GLP [vivone2018full] GLP Fusion Network(ours)
IHS Fusion Network [deng2022machine] Gram Schmidt [dalla2015global]
Figure 10: Fusion output comparisons across 60 m atmospheric bands. The figure show the Low Resolution input, Generalized laplacian pyramid (GLP) baseline method [vivone2018full], our GLP inspired neural network, intensity-hue-saturation (IHS) transform, and Fusion network baseline [deng2022machine] and Gram Schmidt [dalla2015global] method.
Low Resolution GLP [vivone2018full] GLP Fusion Network (ours)
IHS Fusion Network [deng2022machine] Gram Schmidt [dalla2015global]
Figure 11: Fusion output comparisons across 60 m atmospheric bands. The figure show the Low Resolution input, Generalized laplacian pyramid (GLP) baseline method [vivone2018full], our GLP inspired neural network, intensity-hue-saturation (IHS) transform, and Fusion network baseline [deng2022machine] and Gram Schmidt [dalla2015global] method.